How long does it take an investment of $7,024.40 take to turn into $8,364.50 if the appropriate interest rate is 1.6%? You may round your answer to the nearest whole number
Answer #1

FV = PV(1+r)^n

8364.5 = 7024.40 (1+0.016)^n

1.016^n = 8364.50 / 7024.40

= 1.1908

Take Log on Both sides

Log [ 1.016^n ] Log [ 1.1908 ]

Log [ a^b ] = b * log [ a ]

n * Log [ 1.016 ] = Log [ 1.1908 ]

n * 0.0069 = 0.0758

n = 0.0758 / 0.0069

= 11

period Required is 11